OpenStudy (cooltowl):

With 4.95 g of baking soda, how many moles of Na2O should be formed? Using stoichiometric calculations 2 NaHCO3(s) → 1 Na2O(s) + 2 CO2(g) + 1 H2O(g)

OpenStudy (anonymous):

Start with baking soda. \[g NaHCO _{3} \times \frac{ mol NaHCO _{3} }{ gNaHCO _{3} } \times \frac{ mol Na _{2}O }{ mol NaHCO _{3} } \times \frac{ gNa _{2}O }{ mol Na _{2}O }\] Can you fill in the missing numbers?

OpenStudy (aaronq):

you're gonna make half of the moles of sodium bicarbonate, look at their coefficients.